# Does the suppression shown at LHC in O-O collisions follow the systematics obtained for A-A collisions ??

### v1

@@ Abstract
In this letter we want to see to what extent recent experimental results obtained for $\pi^{0}$ suppression in O-O collisions at $\sqrt{s_{NN}}$=5.36 TeV fit into the systematics for much heavier systems. The systematics with which the comparison is made was published a few years ago [Pet_1] in terms of charged particles suppression $R_{AA}$ as a function of $\langle N_{part}\rangle$ and $\langle dN_{ch}/d\eta\rangle$ and the newly introduced $R_{AA}^{N}=\frac{(\frac{d^{2}N}{dp_{T}d\eta}/\langle\frac{dN_{ch}}{d\eta}\rangle)^{cen}}{(\frac{d^{2}N}{dp_{T}d\eta}/\langle\frac{dN_{ch}}{d\eta}\rangle)^{pp,INEL}}$ as a function of $\langle dN_{ch}/d\eta\rangle$. The values of $R_{AA}$ and $R_{AA}^{N}$ at $\langle dN_{ch}/d\eta\rangle$ and $\langle N_{part}\rangle$ experimentally measured and estimated by Glauber MC, respectively, for O-O collisions are in good agreement with the systematics obtained for A-A collisions.
@@ I Introduction
The studies of $p$ collisions at LHC energies up to very high charged particle multiplicities, revealed similarities between $p$ and Pb-Pb in the behaviour of different observables, like near-side long range pseudorapidity correlations [CMS1], the ($\langle\beta_{T}\rangle$ - $T_{kin}^{fo}$) correlation as a function of charged particle multiplicity [Cristi1], azimuthal angular correlations [CMS2], geometrical scaling [Pet_2, Pet_3, Pet_4], etc. Based on the suppression studies in Cu-Cu and Au-Au collisions at RHIC and Xe-Xe and Pb-Pb collisions at LHC and the dependence of $R_{AA}^{N}=\frac{(\frac{d^{2}N}{dp_{T}d\eta}/\langle\frac{dN_{ch}}{d\eta}\rangle)^{cen}}{(\frac{d^{2}N}{dp_{T}d\eta}/\langle\frac{dN_{ch}}{d\eta}\rangle)^{pp,INEL}}$ on $\langle dN_{ch}/d\eta\rangle$, in ref. [Pet_1], it was argued why even in high charged particle multiplicity events in $p$ collisions at LHC, in the limit of current experimental uncertainties, no suppression was observed, although similarities to Pb-Pb collisions for other observables have been highlighted. Recently, at the LHC energy of $\sqrt{s_{NN}}$=5.36 TeV, O-O collsions were measured and preliminary $R_{AA}$ results were reported by ALICE [Stra, Mar] and CMS [CMSS] Collaborations for $\pi^{0}$ and charged particles, respectively. In this letter we are comparing the ALICE results with the results of a systematic study of $R_{AA}$ and $R_{AA}^{N}$ as a function of $\langle dN_{ch}/d\eta\rangle$ and $\langle N_{part}\rangle$ for A-A collisions.
@@
In the paper cited above [Pet_1] was also presented a dependence on $\langle N_{part}\rangle$ of $\langle N_{bin}\rangle$$\langle dN_{ch}/d\eta\rangle^{A-A}/\langle dN_{ch}/d\eta\rangle^{pp}$] for Cu-Cu and Au-Au at $\sqrt{s_{NN}}$=200 GeV, Xe-Xe at $\sqrt{s_{NN}}$=5.44 TeV, Pb-Pb at $\sqrt{s_{NN}}$=2.76 TeV and 5.02 TeV using for $\langle dN_{ch}/d\eta\rangle$ the values reported in [STAR_dNchdeta, BES_CuCu_dNchdeta, XeXe_data, PbPb_276_dNchdeta, PbPb_502_dNchdeta], $N_{bin}>$ estimated within the Glauber MC approach [Loiz_1] and the $\langle\frac{dN_{ch}}{d\eta}\rangle)^{pp}$ INEL values at the corresponding collision energies (Fig.1). Within the error bars, a good scaling is observed. The rather large values of the error bars are due to uncertainty in $\langle N_{bin}\rangle$ estimates [Loiz_1]. In Fig.1 are represented by yellow squares the values corresponding to O-O collisions at $\sqrt{s_{NN}}$=5.36 TeV for different centralities, using $\langle dN_{ch}/d\eta\rangle$ reported by the ALICE [Stra, Mar] Collaboration. $\langle N_{part}\rangle$ and $\langle N_{bin}\rangle$ estimated within the Glauber MC approach were taken from [Loiz] for the corresponding centralities. As can be observed, the values for O-O follow the systematics obtained in heavy ion collisions.
Figure 1: $\langle dN_{bin}\rangle$$\langle dN_{ch}/d\eta\rangle^{A-A}/\langle dN_{ch}/d\eta\rangle^{pp}$] values in O-O collision at $\sqrt{s_{NN}}$=5.36 TeV are presented by yellow squares using recently measured $\langle dN_{ch}/d\eta\rangle$ for different centralities by the ALICE [Stra, Mar] Collaboration on top of Fig.14 from [Pet_1]
Figure 2: $R_{AA}$ as a function of $\langle N_{part}\rangle$ from [Pet_2] over which the corresponding experimental value for O-O is represented by the yellow box.
@@ III $R_{AA}$ - $\langle N_{part}\rangle$ dependence
As it was presented in a previous paper [Pet_2], the transverse overlap area of the nucleons suffering more that single collisions $S_{\perp}^{core}$ shows a very good scaling as a function of $N_{part}$, both being estimated within the Glauber MC approach for the Au-Au at energies measured at RHIC and Pb-Pb at energies measured at LHC. Similar results were obtained for $S_{\perp}^{var,core}$ for Cu-Cu, Au-Au at top energy at RHIC and Xe-Xe and Pb-Pb at energies measured at LHC. The aspect ratios of the transverse overlap area $S_{\perp}$ for different systems and energies are also rather similar. With the assumption that the average path length of a parton traveling the deconfined zone $L^{2}\propto S_{\perp}$ one could expect a good scaling of the average suppression as a function of $N_{part}$ at a given collision energy. This was presented in ref. [Pet_1]. For O-O at $\sqrt{s_{NN}}$=5.36 TeV, an estimate of $N_{part}$=10.8 for 0-100% centrality based on the Glauber MC approach was reported in a recent publication [Loiz]. In Fig.2 by the yellow box is represented the $R_{AA}$ value reported as preliminary results by the ALICE [Stra, Mar] Collaboration at the corresponding $N_{part}$ on top of the compilation presented in [Pet_1]. The agreement between the O-O experimental result and the $N_{part}$ dependence of $R_{AA}$ obtained in Xe-Xe and Pb-Pb at similar collision energies is good.
@@ IV $R_{AA}-\langle dN_{ch}/d\eta\rangle$ dependence
Within the phenomenological parton energy loss approach [Xu], the fractional energy loss is:
@@
where a and b take different values [Xu]. This could explain the collision energy dependence of $R_{AA}$ as a function of $N_{part}$, the temperature being rather different as far as $\langle dN_{ch}/d\eta\rangle$ is different for a given value of $N_{part}$ [Pet_1]. Based on this argument, in [Pet_1] we also presented $R_{AA}$ as a function of $dN_{ch}/d\eta$ for the same systems. In such a representation the energy dependence of the scaling does not appear anymore although the aspect ratio of $S_{\perp}$ for different systems and collision energy is rather different. The study of the azimuthal dependence of $R_{AA}$ done by the PHENIX Collaboration for Au-Au collision at $\sqrt{s_{NN}}$=200 GeV and 30-40% centrality showed a larger suppression out of plane relative to the one in-plane [PHE]. It seems that this effect is less important in the average value of $R_{AA}$ and the temperature dependence of the energy loss in the deconfined matter plays the main role.
Figure 3: $R_{AA}$ as a function of $\langle dN_{ch}/d\eta\rangle$ from [Pet_2] on which the corresponding experimental value for O-O is represented by the yellow box.
Taking the measured value of $\langle dN_{ch}/d\eta\rangle$ for O-O collisions at $\sqrt{s_{NN}}$=5.36 TeV reported by the ALICE [Stra, Mar] Collaboration, the result is presented in Fig.3 by the yellow square on top of the compilation presented in Fig.8 of reference [Pet_2]. The $R_{AA}$ for the O-O case is in good agreement with the results obtained for much heavier systems at RHIC and LHC at the same $\langle dN_{ch}/d\eta\rangle$.
V $R_{AA}^{N}-\langle dN_{ch}/d\eta\rangle$ dependence
In the same paper [Pet_1] we also presented, as a function of $\langle dN_{ch}/d\eta\rangle$, the systematics of $R_{AA}^{N}$, a model independent quantity, obtained as the ratio of the $p_{T}$ spectra in A-A collisions to the one in INEL $p$ collisions at the same energy, each of them normalised to the corresponding charged particle densities, for all available centralities in A-A collisions.
$R_{AA}^{N}=\frac{(\frac{d^{2}N}{dp_{T}d\eta}/\langle\frac{dN_{ch}}{d\eta}\rangle)^{cen}}{(\frac{d^{2}N}{dp_{T}d\eta}/\langle\frac{dN_{ch}}{d\eta}\rangle)^{pp,INEL}}$ (2)
In Fig.4 $R_{AA}^{N(O-O)}$ at $\sqrt{s_{NN}}$=5.36 TeV and $R_{AA}^{N(Pb-Pb)}$ at $\sqrt{s_{NN}}$=2.76 TeV and 5.02 TeV at the same $\langle dN_{ch}/d\eta\rangle$ show a rather good agreement in the limit of the present uncertainties of the analysis.
Figure 4: $R_{AA}^{N}$ as a function of $\langle dN_{ch}/d\eta\rangle$ from [Pet_2] over which the corresponding experimental value for O-O is represented by the yellow box.
VI Conclusions
Although the alpha-like cluster structure of the colliding O nuclei could result in peculiar behaviour of many observables, it seems that the average value of $R_{AA}$ is in very good agreement with those extracted from Pb-Pb and Xe-Xe collisions at a similar collision energy at the same $\langle dN_{ch}/d\eta\rangle$ or $\langle N_{part}\rangle$ values. The expected results from detailed data analysis of the information collected in light heavy ion collisions at LHC will show to which extent the previously established systematics in heavy ion collisions and $p$ at LHC energies [Pet_1, Pet_3, Pet_4, Top] are also followed in the case of light heavy ion collisions.

### v2

@@ Abstract
In this paper we present to what extent recent experimental results obtained for $\pi^{0}$ suppression in O-O collisions at $\sqrt{s_{NN}}$=5.36 TeV fit into the systematics for much heavier systems. The systematics with which the comparison is made was published a few years ago [14] in terms of charged particles suppression $R_{AA}$ as a function of $\langle N_{part}\rangle$ and $\langle dN_{ch}/d\eta\rangle$. The values of $R_{AA}$ at $\langle dN_{ch}/d\eta\rangle$ and $\langle N_{part}\rangle$ experimentally measured and estimated by Glauber MC, respectively, for O-O collisions are in good agreement with the systematics obtained for A-A collisions.
@@ I Introduction
The studies of $p$ collisions at LHC energies up to very high charged particle multiplicities, revealed similarities between $p$ and Pb-Pb in the behaviour of different observables, like near-side long range pseudorapidity correlations [8], the ($\langle\beta_{T}\rangle$ - $T_{kin}^{fo}$) correlation as a function of charged particle multiplicity [6], azimuthal angular correlations [9], geometrical scaling [13, 15, 16], etc. Based on the suppression studies in Cu-Cu and Au-Au collisions at RHIC and Xe-Xe and Pb-Pb collisions at LHC and the dependence of $R_{AA}^{N}=\frac{(\frac{d^{2}N}{dp_{T}d\eta}/\langle\frac{dN_{ch}}{d\eta}\rangle)^{cen}}{(\frac{d^{2}N}{dp_{T}d\eta}/\langle\frac{dN_{ch}}{d\eta}\rangle)^{pp,INEL}}$ on $\langle dN_{ch}/d\eta\rangle$ [14], it was argued why even in high charged particle multiplicity events in $p$ collisions at LHC, in the limit of current experimental uncertainties, no suppression was observed, although similarities to Pb-Pb collisions for other observables have been highlighted. Recently, at the LHC energy of $\sqrt{s_{NN}}$=5.36 TeV, O-O collsions were measured and preliminary $R_{AA}$ results were reported by ALICE [18, 12] and CMS [7] Collaborations for $\pi^{0}$ and charged particles, respectively. In this letter we are comparing the ALICE results with the results of a systematic study of $R_{AA}$ as a function of $\langle dN_{ch}/d\eta\rangle$ and $\langle N_{part}\rangle$ for A-A collisions.
@@
In the paper cited above [14] was also presented a dependence on $\langle N_{part}\rangle$ of $\langle N_{bin}\rangle$$\langle dN_{ch}/d\eta\rangle^{A-A}/\langle dN_{ch}/d\eta\rangle^{pp}$] for Cu-Cu and Au-Au at $\sqrt{s_{NN}}$=200 GeV, Xe-Xe at $\sqrt{s_{NN}}$=5.44 TeV, Pb-Pb at $\sqrt{s_{NN}}$=2.76 TeV and 5.02 TeV using for $\langle dN_{ch}/d\eta\rangle$ the values reported in [2, 5, 3, 1, 4], $N_{bin}>$ estimated within the Glauber MC approach [10] and the $\langle\frac{dN_{ch}}{d\eta}\rangle)^{pp}$ INEL values at the corresponding collision energies (Fig.1). Within the error bars, a good scaling is observed. The rather large values of the error bars are due to uncertainty in $\langle N_{bin}\rangle$ estimates [10]. In Fig.1 are represented by yellow squares the values corresponding to O-O collisions at $\sqrt{s_{NN}}$=5.36 TeV for different centralities, using $\langle dN_{ch}/d\eta\rangle$ reported by the ALICE Collaboration [18, 12]. $\langle N_{part}\rangle$ and $\langle N_{bin}\rangle$ estimated within the Glauber MC approach were taken from [11] for the corresponding centralities. As can be observed, the values for O-O collisions follow the systematics obtained in heavy ion collisions.
Figure 1: $\langle dN_{bin}\rangle$$\langle dN_{ch}/d\eta\rangle^{A-A}/\langle dN_{ch}/d\eta\rangle^{pp}$] values in O-O collisions at $\sqrt{s_{NN}}$=5.36 TeV are presented by yellow squares using recently measured $\langle dN_{ch}/d\eta\rangle$ for different centralities by the ALICE Collaboration [18, 12] on top of Fig.14 from [14]
Figure 2: $R_{AA}$ as a function of $\langle N_{part}\rangle$ from [13] over which the corresponding experimental value for O-O is represented by the yellow box.
@@ III $R_{AA}$ - $\langle N_{part}\rangle$ dependence
As it was presented in a previous paper [13], the transverse overlap area of the nucleons suffering more that single collisions $S_{\perp}^{core}$ shows a very good scaling as a function of $N_{part}$, both being estimated within the Glauber MC approach for the Au-Au at energies measured at RHIC and Pb-Pb at energies measured at LHC. Similar results were obtained for $S_{\perp}^{var,core}$ for Cu-Cu, Au-Au at top energy at RHIC and Xe-Xe and Pb-Pb at energies measured at LHC. The aspect ratios of the transverse overlap area $S_{\perp}$ for different systems and energies are also rather similar. With the assumption that the average path length of a parton traveling the deconfined zone $L^{2}\propto S_{\perp}$ one could expect a good scaling of the average suppression as a function of $N_{part}$ at a given collision energy. This was presented in ref. [14]. For O-O at $\sqrt{s_{NN}}$=5.36 TeV, an estimate of $N_{part}$=10.8 for 0-100% centrality based on the Glauber MC approach was reported in a recent publication [11]. In Fig.2 by the yellow box is represented the $R_{AA}$ value reported as preliminary results by the ALICE Collaboration [18, 12] at the corresponding $N_{part}$ on top of the compilation presented in [14]. The agreement between the O-O experimental result and the $N_{part}$ dependence of $R_{AA}$ obtained in Xe-Xe and Pb-Pb at similar collision energies is very good.
@@ IV $R_{AA}-\langle dN_{ch}/d\eta\rangle$ dependence
Within the phenomenological parton energy loss approach, the fractional energy loss is given by:
@@
where a and b take different values [20]. This could explain the collision energy dependence of $R_{AA}$ as a function of $N_{part}$, the temperature being rather different as far as $\langle dN_{ch}/d\eta\rangle$ is different for a given value of $N_{part}$ [14]. Based on this argument, in [14] we also presented $R_{AA}$ as a function of $dN_{ch}/d\eta$ for the same systems. In such a representation the energy dependence of the scaling does not appear anymore although the aspect ratio of $S_{\perp}$ for different systems and collision energy is rather different. The study of the azimuthal dependence of $R_{AA}$ done by the PHENIX Collaboration for Au-Au collision at $\sqrt{s_{NN}}$=200 GeV and 30-40% centrality showed a larger suppression out of the reaction plane relative to the one in-plane [17]. It seems that this effect is less important in the average value of $R_{AA}$ and the temperature dependence of the energy loss in the deconfined matter plays the main role.
Figure 3: $R_{AA}$ as a function of $\langle dN_{ch}/d\eta\rangle$ from [13] on which the corresponding experimental value for O-O is represented by the yellow box.
Taking the measured value of $\langle dN_{ch}/d\eta\rangle$ for O-O collisions at $\sqrt{s_{NN}}$=5.36 TeV reported by the ALICE Collaboration [18, 12] , the result is presented in Fig.3 by the yellow square on top of the compilation presented in Fig.8 of reference [13]. The $R_{AA}$ for the O-O case is in good agreement with the results obtained for much heavier systems at RHIC and LHC at the same $\langle dN_{ch}/d\eta\rangle$.
V Conclusions
Although the alpha-like cluster structure of the colliding O nuclei could result in peculiar behaviour of many observables, it seems that the average value of $R_{AA}$ is in very good agreement with those extracted from Pb-Pb and Xe-Xe collisions at a similar collision energy at the same $\langle dN_{ch}/d\eta\rangle$ or $\langle N_{part}\rangle$ values. The expected results from detailed multi differetial data analysis of the information collected in alpha-like light ion collisions at LHC will show to which extent the previously established systematics in heavy ion collisions and $p$ at LHC energies [14, 15, 16, 19] are also followed in the case of such collisions.